# PAD: Phase-Amplitude Decoupling Fusion for Multi-Modal Land Cover Classification

### v1

@@ IV-C1 Evaluations on WHU-OPT-SAR
IV-C2 Evaluations on DDHR-SK
As shown in Table III, PAD exhibits superior boundary segmentation performance on the relatively simplified scenes of DDHR-SK. Compared with leading unimodal models (e.g., Mask2former-RGB [18], achieving 90.25% mIoU and 94.76% mF1, DANet-SAR [13], yielding 83.31% mIoU and 90.28% mF1) and the multi-modal baseline MRFS [25] (93.57% mIoU, 82.24% mKappa), PAD attains new state-of-the-art performance with 94.59% mIoU, 97.65% OA, 82.24% mKappa, and 96.60% mF1. For boundary-sensitive categories, PAD achieves 99.47% IoU on Water and 86.12% IoU on Road edges, surpassing MRFS by 2.44% and 0.08%, respectively.
Figure 7: Visualization Results on DDHR-SK. Four boundary-focused cases are shown: (a) coastal demarcation, (b) water–cropland interface, (c) vegetation–building complex, and (d) road network topology. PAD is compared with leading unimodal methods (Mask2former-RGB, DANet-SAR) and state-of-the-art multi-modal approaches (CDDFuse, CMX, MRFS). For each case, the panels display the RGB input, SAR coherence map, pseudo-color ground truth, and prediction maps from six methods. Critical boundary regions are highlighted with orange boxes; further analysis is provided in Section IV-C.
The visual results further reveal PAD’s capability in cross-modal boundary refinement. In Fig. 7(a), PAD accurately reconstructs periodically submerged coastal lines via SAR coherence analysis, whereas RGB models tend to confuse tidal sediments with terrestrial areas. In Fig. 7(b), PAD leverages SAR features to delineate water infiltration boundaries in irrigated fields, overcoming the over-segmentation issues common to optical sensors. Moreover, in Fig. 7(c) and Fig. 7(d), PAD achieves continuous road networks under vegetation occlusion and precise material boundaries in vegetation–building complexes, respectively.
IV-C3 Ablation Study
We performed comprehensive ablation studies on WHU-OPT-SAR to validate the contributions of the core components of our approach: Amplitude Spectrum Fusion module (ASF), Phase Spectrum Correction module (PSC), and Multi-Stage Fusion (MSF).
@@ IV-C3 Ablation Study
Eight model variants (see Table IV) were tested. The baseline variant (I), which replaces MSF with early fusion and substitutes the PAD (ASF+PSC) with pixel-wise additive fusion, achieves the lowest mIoU (54.98%). Enabling MSF alone (II) improves mIoU by 0.39%, while using ASF alone (III) attains 55.86% mIoU, and solely PSC (IV) adds a smaller gain. Combining MSF with either ASF (V) or PSC (VI) yields similar improvements (mIoU around 56.11–56.12%). Integrating all components achieves the best performance (56.26% mIoU), demonstrating the effective synergy of all modules.
TABLE V: Ablation Experiment Results (%) on Loss Function Weights on WHU-OPT-SAR.
Configuration IoU per Category aAcc Kappa mF1 mIoU
@@
We conducted weight ablation studies (see Table V) to examine the impact of the trade-off between the amplitude consistency loss ($\mathcal{L}_{\mathrm{amp}}$) and the segmentation loss ($\mathcal{L}_{\mathrm{seg}}$). The best performance (mIoU = 56.26%) is achieved with $\lambda_{1}=0.4$ and $\lambda_{2}=0.1$. Increasing $\lambda_{1}$ above 0.4 degrades performance, notably in under-represented classes, while an overly low $\lambda_{2}$ significantly diminishes performance, confirming that segmentation supervision must remain the dominant objective.
TABLE VI: Ablation Results (%) on the Initial Frequency Radius on WHU-OPT-SAR.
Configuration IoU per Category aAcc Kappa mF1 mIoU
@@
The initial frequency radius parameter $r$, which controls the proportion of preserved low-frequency components in our phase-aware filtering, was varied (see Table VI). Our experiments show that $r=0.1$ achieves an optimal balance (56.26% mIoU). Values that overly suppress low-frequency information or over-smooth the features lead to inferior performance, though the method remains robust for $r\in[0.1,10]$.
@@ IV-D Visualization at Distribution Level
To further analyze the feature distribution, we extracted multi-scale features from the RGB and SAR encoders, as well as from the fusion modules. After upsampling to the original resolution, the features were concatenated channel-wise. To balance computational efficiency with preserving discriminative information, a two-stage dimensionality reduction was applied: 1) random spatial sampling, followed by 2) principal component analysis (PCA), retaining 50 principal components (accounting for over 89% of the variance). The final visualization was generated using t-SNE [54] (with a perplexity of 30 and 1,000 iterations).
@@
Figure 8: t-SNE Visualization of Cross-modal Feature Distribution Evolution. (a) RGB encoder features exhibit spectral continuity but have blurred boundaries; (b) SAR encoder features, while strongly sensitive to geometry, are affected by speckle noise; (c) PAD-fused features display enhanced intra-class compactness and improved semantic separation.
As illustrated in Fig. 8, the PAD-fused features exhibit enhanced intra-class compactness and improved inter-class separability. Quantitatively, the mean intra-class standard variance is reduced by 20.39% for Road, 13.16% for Water, 6.46% for Forest, and 4.43% for Farmland, compared to the best unimodal baselines. In addition, significant increases in Euclidean distances between critical class pairs, such as a 75.75% increase in the City/Village pair (66.56 to 116.95) and a 13.29% increase for the Farmland/Forest pair (107.64 to 121.89), indicate that the fused features capture cross-modal discriminative patterns effectively.
Overall, these experimental results confirm that the PAD framework significantly improves classification accuracy and prediction consistency by effectively integrating multi-modal information and suppressing noise, while maintaining the discriminative spectral and geometric characteristics inherent in the data.
@@ V Conclusion
In this paper, we proposed the Phase-Amplitude Decoupling (PAD) framework for fusing SAR and RGB modalities in land cover classification. Our frequency-domain analysis revealed that phase components are consistent across modalities, while amplitude spectra capture modality-specific details, providing the first quantitative physical prior for multi-source remote sensing fusion. PAD separates and aligns phase features while adaptively fusing amplitude components. The Phase Spectrum Correction (PSC) module resolves structural discrepancies, and the Amplitude Spectrum Fusion (ASF) module dynamically integrates SAR’s high-frequency details with RGB’s low-frequency context. Extensive experiments demonstrate that PAD outperforms existing methods in both accuracy and robustness. Future work will focus on improving phase alignment, particularly under significant geometric distortions due to imprecise cross-modal registration.

### v2

@@ IV-C4 Ablation Study
Eight variants (see Table V) were evaluated. The baseline (I), which replaces MSF with early fusion and substitutes PAD (ASF+PSC) with pixel-wise addition, achieves the lowest mIoU (54.98%). Enabling MSF alone (II) yields a +0.39% gain, ASF alone (III) reaches 55.86% mIoU, and PSC alone (IV) provides a smaller improvement. Combining MSF with ASF (V) or PSC (VI) yields similar gains (56.11-56.12% mIoU). Integrating all three modules achieves the best performance (56.26% mIoU), demonstrating their effective synergy.
TABLE VI: Ablation Study on Fusion Modules (%) for WHU-OPT-SAR. We compare alternative amplitude-only, phase-only, and full fusion designs by integrating each into our framework.
Fusion Module IoU per Category OA mKappa mF1 mIoU
Farmland City Village Water Forest Road Others
Amplitude FelNFN [16] 66.93 55.80 45.15 61.27 82.77 35.54 19.17 82.95 47.31 66.46 52.38
FCENet [14] 68.27 55.18 46.79 61.17 83.43 37.09 16.71 83.43 49.82 66.52 52.66
FGSwin [53] 69.87 57.43 49.03 65.04 84.12 43.16 22.26 84.51 53.53 69.70 55.85
Ours 69.91 58.51 49.53 65.52 84.05 42.63 22.58 84.55 54.53 69.93 56.11
Phase FelNFN [16] 60.66 53.25 31.99 54.14 80.64 19.84 15.02 79.42 33.32 58.89 45.08
Channel Attention 69.58 58.66 49.78 64.28 83.96 43.43 22.65 84.39 53.54 69.93 56.05
Ours 69.98 57.32 49.36 65.62 84.12 43.42 23.01 84.57 53.80 69.98 56.12
Whole FelNFN [16] 66.43 51.97 44.36 58.48 82.80 32.33 12.64 82.33 47.35 63.63 49.86
FCENet++ [14] 66.87 53.01 46.41 60.30 82.64 30.21 14.51 82.78 43.83 64.33 50.56
FGSwin [53] 69.88 57.86 49.48 65.26 84.15 42.49 22.36 84.56 53.41 69.76 55.92
Frefusion [15] 69.91 57.99 49.35 64.22 84.00 42.72 22.46 84.41 53.98 69.69 55.81
SFINet++ [46] 68.23 55.65 47.31 62.35 83.10 38.48 19.04 83.50 49.56 67.43 53.45
Ours 69.85 58.09 49.71 65.39 84.05 43.24 23.47 84.56 53.96 70.14 56.26
Fusion Modules
To further validate ASF and PSC, we replaced each with five existing frequency-domain designs, FelNFN [16], FCENet [14], FGSwin [53], Frefusion [15], and SFINet++ [46], while keeping the rest of our framework unchanged (see Table VI). These methods represent common strategies such as fixed-kernel spectral convolution (FelNFN), frequency-spatial interaction via Hadamard products/matrix multiplication (FCENet), learned spectral filtering (FGSwin), attention-based cross-domain fusion (Frefusion), and amplitude-phase parallel processing (SFINet++). Our ASF outperforms other amplitude-specific methods (56.11% vs. 55.85% mIoU for FGSwin), and PSC consistently exceeds phase baselines (56.12% vs. 56.05% for channel attention and 45.08% for FelNFN). At the full-fusion level, our integrated design still leads (56.26% mIoU), indicating both modules are effective in isolation and synergistic when combined. These results confirm that adaptive frequency modulation (ASF) and spatially guided phase correction (PSC) offer structurally unique and more effective alternatives to existing paradigms.
TABLE VII: Ablation Study on Loss Function Weights (%) for WHU-OPT-SAR. We vary $\lambda_{1}$ with fixed $\lambda_{2}=0.1$ and $\lambda_{2}$ with fixed $\lambda_{1}=0.4$.
Configuration IoU per Category OA mKappa mF1 mIoU
@@
We conducted weight ablations (see Table VII) to assess the trade-off between the amplitude consistency loss $\mathcal{L}_{\mathrm{amp}}$ and segmentation loss $\mathcal{L}_{\mathrm{seg}}$. The best mIoU (56.26%) is achieved with $\lambda_{1}=0.4$ and $\lambda_{2}=0.1$. Increasing $\lambda_{1}$ above 0.4 degrades performance, especially on under-represented classes, while too small $\lambda_{2}$ significantly reduces accuracy, confirming segmentation supervision must remain dominant.
TABLE VIII: Ablation Study on Initial Frequency Radius (%) for WHU-OPT-SAR. We vary the trainable radius parameter in the ASF module.
Radius IoU per Category OA mKappa mF1 mIoU
@@
We varied the initial frequency radius $r$, which controls the preserved low-frequency proportion in ASF (see Table VIII). The optimal balance is at $r=0.1$ (56.26% mIoU). Values that overly suppress or over-smooth low-frequency content degrade performance, although the method remains robust for $r\in[0.1,10]$.
@@ IV-D Visualization at Distribution Level
To further analyze feature distributions, we extracted multi-scale features from the RGB and SAR encoders and from the fusion modules. After upsampling to the original resolution, we concatenated features channel-wise. To balance computational efficiency with discriminative power, we applied a two-stage dimensionality reduction: a) random spatial sampling, and b) principal component analysis (PCA), retaining 50 components (over 89% variance). The final visualization was generated using t-SNE [54] (perplexity = 30, 1,000 iterations).
@@
Figure 8: t-SNE Visualization of Cross-modal Feature Distribution Evolution. (a) RGB encoder features show spectral continuity but blurred class boundaries. (b) SAR encoder features capture geometric details but are affected by speckle noise. (c) PAD-fused features exhibit enhanced intra-class compactness and improved inter-class separation.
As shown in Fig. 8, PAD-fused features exhibit enhanced intra-class compactness and improved inter-class separability. Quantitatively, mean intra-class standard variance is reduced by 20.39% for roads, 13.16% for water, 6.46% for forest, and 4.43% for farmland, compared to the best unimodal baselines. Moreover, Euclidean distances between critical class pairs increase significantly: City/Village by 75.75% (66.56 $to$ 116.95) and Farmland/Forest by 13.29% (107.64 $to$ 121.89), indicating that PAD captures cross-modal discriminative patterns effectively.
TABLE IX: Statistical Comparison Between Models w/ and w/o ASF.
Metric w/ ASF w/o ASF
Mean Pearson Correlation 0.0364 0.0369
Mean Spearman Correlation 0.0482 0.0488
Paired t-test p-value 0.0016
Wilcoxon test p-value $6.4\times 10^{-7}$
Cohen’s $d$ (effect size) -0.038
IV-E Noise Sensitivity of ASF
To assess whether ASF amplifies high-frequency speckle noise in SAR images, we conducted a correlation analysis using the Equivalent Number of Looks (ENL), a standard measure of speckle. For each image, we computed pixel-wise ENL maps and then measured the Pearson and Spearman correlations between ENL and classification accuracy. As shown in Table IX, the model with ASF exhibits nearly identical correlation coefficients to the baseline, differences in mean correlation are below 0.001, and Cohen’s $d=-0.038$ indicates a negligible effect size. These results confirm that ASF does not increase noise sensitivity and remains robust in low-SNR regions.
Overall, these results confirm that PAD significantly improves classification accuracy and prediction consistency by integrating multi-modal information, suppressing noise, and preserving discriminative spectral and geometric characteristics inherent in the data.
@@ V Conclusion
In this paper, we proposed the Phase-Amplitude Decoupling (PAD) framework for fusing SAR and RGB modalities in land cover classification. Our frequency-domain analysis quantitatively demonstrates that phase components remain consistent across modalities, while amplitude spectra capture modality-specific details, providing the first physical prior for multi-source remote sensing fusion. PAD disentangles and aligns phase features via the Phase Spectrum Correction (PSC) module, and adaptively fuses amplitude components using the Amplitude Spectrum Fusion (ASF) module, which integrates SAR’s high-frequency details with RGB’s low-frequency context. Extensive experiments on two challenging datasets show that PAD outperforms existing methods in both accuracy and robustness.
Despite its demonstrated effectiveness, PAD assumes well-aligned input modalities and relies on hard-label supervision. In future work, we will integrate self-supervised cross-modal registration refinement and soft-label modeling, and investigate transformer-based frequency fusion to further enhance global context awareness.